OpenStudy (anonymous):

a rectangle is twice as long as it is wide . if its length is decreased by 4 and its width is decresed by 2, its area is decreased by 24. find its original diemensions

OpenStudy (ajprincess):

 Area A=length*width=2w*w New area A1=(2w-4)(w-2) A-24=A1 2w*w-24=(2w-4)(w-2) solve this to find w. then use l=2w to find l. Can u do?